# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 04 Mar 2026 v2

@@ 17.1 Topological Barriers and Infinite-Dimensional Picard-Lefschetz Theory
By analytically continuing the twistor path integral into the complexified phase space, the convergence of the quantum amplitude is governed by middle-dimensional integration cycles, or Lefschetz thimbles [5]. As formulated by Witten in the analytic continuation of Chern-Simons theory [55], the topology of these thimbles—and their intersection pairings—changes strictly when the system crosses a Stokes line, resulting in a discontinuous jump in the physical vacuum state. In our twistor string framework, non-perturbative Stokes phenomena are rigorously suppressed by the conformal boundary geometry. If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour. Exactly at the moment of crossing, these two singularities inevitably collide directly on the physical integration contour $|z|=1$. As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity. The integration loop becomes trapped between the colliding singularities at $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them. This generates a severe pinch singularity, creating an infinite action barrier. In the language of Picard-Lefschetz theory, this impenetrable boundary dynamically locks the Twistor String path integral into a strictly conserved topological sector defined by the enclosed twistor pole number $n$, forbidding unphysical Stokes jumps.
@@ 17.2 Complex Langevin, Cauchy’s Shield, and Flat Valleys
The resolution lies in the analytic continuation of the twistor fields. The CLE directly evolves the boundary data $\mathcal{V}_{bound}(\tau)=\mathcal{V}(z,\tau)\big|_{|z|=1}$. However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk. If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{k}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
$\lambda(\theta)=\sum_{k}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (17.1)
@@
$\lambda(\theta,\tau)=\sum_{k}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (17.2)
@@ 18 Conclusion: The Solution of Planar QCD
Our second main outcome is dynamical: we demonstrated that the internal Majorana fermions (the “elves”) on the worldsheet provide the precise algebraic mechanism required for planar factorization at intersections. Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions. The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/48-1/48 and $23/48$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
In a profound sense, this construction explicitly realizes Edward Witten’s original “Master Field” hypothesis—the existence of a purely classical description for the large-$N_{c}$ limit of QCD. While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, we have discovered its exact geometric realization in loop space. The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface. Consequently, the phase-space path integral over quark loops, once transformed into two-dimensional boundary twistor variables, exhibits exact geometric localization governed by Catastrophe Theory. The true Master Field of Planar QCD is thus a critical classical trajectory in twistor space—a geometric structure conceptually parallel to the spectral curves of Seiberg-Witten theory.
@@
Ultimately, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry. The extraction of the exact mass spectrum reduces from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem. By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition. For the fundamental one-pole trajectory, this yields the unified linear Regge spectrum:
$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (18.1)
where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory. This analytic solution passes flawlessly through experimental data on Figure 15. The only outlier ($3\sigma$) is $K_{3}(2320)$ which is marked as "unconfirmed" in the PDG tables. The exact mass spectrum of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
@@ Acknowledgements
No data was created in this paper. The analytic computations of the Magnus expanstion of the Momentum loop equations were performed in the Mathematica ® notebook published on the Wolfram Cloud[37].
@@ D.2 The Euclidean Metric and Parity Projection
To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$. The flat integration measure transforms as $dzd\bar{z}/|z|^{2}=d\sigma_{E}^{2}+d\tau_{E}^{2}$. Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
@@
$ds^{2}_{E}=\exp\left(2\rho\right)(d\sigma_{E}^{2}+d\tau_{E}^{2})$
$=R^{2}\left(1+e^{-2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
To evaluate the physical states, we analytically continue this metric into Minkowski spacetime via a double Wick rotation. The angular parameter $\tau_{E}$ serves as the target space time. We perform a Wick rotation on the spatial worldsheet coordinate into the physical light-cone regime:
@@
Under this transformation, the exponential strictly rotates into a complex phase: $e^{-2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})-\mathrm{i}\sin(2\sigma_{M})$.
Because the macroscopic QCD string is strictly parity-conserving, the physical state is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations. The conjugate right-chiral surface identically provides the inverse complex phase $e^{+2\mathrm{i}\sigma_{M}}$. Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the metric:
$\text{Re}\left[\exp\left(2\rho\right)\right]=R^{2}\big(1+\cos(2\sigma_{M})\big)=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact physical Minkowski metric is:
$ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (D.9)
To explicitly verify that this describes the canonical rotating string, we change from the conformal spatial coordinate $\sigma_{M}$ to the physical target space radial coordinate $r$. For a rigidly rotating string, the fractional radius is $r=\sin(\sigma_{M})$. The differentials transform as:
$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (D.10)
Simultaneously, the time component of the metric incorporates the trigonometric identity $\cos^{2}(\sigma_{M})=1-\sin^{2}(\sigma_{M})=1-r^{2}$. Substituting these directly into the Minkowski metric, we obtain:
$ds^{2}_{M}=2R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (D.11)
This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})d\tau_{E}^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light, but exactly halved in overall magnitude relative to the generic three-dimensional configuration.
@@ Appendix E The Pinch Singularity and the Topological Barrier
To mathematically illustrate the appearance of the conjugate pole at $1/\bar{z}_{0}$ and the resulting impenetrable topological barrier at the unit circle, consider the evaluation of the Liouville kinetic anomaly term in the effective action. In for a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
$I=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
We can rewrite the integrand as a total derivative with respect to $\bar{z}$:
$I\propto\int_{D}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $|z|=1$:
$I\propto\oint_{|z|=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
@@
$I\propto\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
@@ Appendix F Algebraic Derivation of Spin Holonomy and Trace T
$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{4\pi i}&0\\ 0&e^{-4\pi i}\end{pmatrix}$ (F.4)
@@
The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulkates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$. which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\mathrm{i}\pi n$ depending on the parity of $L$.

### 12 Mar 2026 v3

@@ 16.1 Topological Barriers and Infinite-Dimensional Picard-Lefschetz Theory
By analytically continuing the twistor path integral into the complexified phase space, the convergence of the quantum amplitude is governed by middle-dimensional integration cycles, or Lefschetz thimbles [4]. As formulated by Witten in the analytic continuation of Chern-Simons theory [55], the topology of these thimbles—and their intersection pairings—changes strictly when the system crosses a Stokes line, resulting in a discontinuous jump in the physical vacuum state. In our twistor string framework, non-perturbative Stokes phenomena are rigorously suppressed by the conformal boundary geometry. If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour. Exactly at the moment of crossing, these two singularities collide directly on the physical integration contour $|z|=1$. As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity. The integration loop becomes trapped between the colliding singularities at $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them. This generates a severe pinch singularity, creating an infinite action barrier. In the language of Picard-Lefschetz theory, this impenetrable boundary dynamically locks the Twistor String path integral into a strictly conserved topological sector defined by the enclosed twistor pole number $n$, forbidding unphysical Stokes jumps.
@@ 16.2 Complex Langevin, Cauchy’s Shield, and Flat Valleys
The resolution lies in the analytic continuation of the twistor fields. The CLE directly evolves the boundary data $\mathcal{V}_{bound}(\tau)=\mathcal{V}(z,\tau)\big|_{|z|=1}$. However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk. If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{i}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
$\lambda(\theta)=\sum_{i}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (16.1)
@@
$\lambda(\theta,\tau)=\sum_{i}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (16.2)
@@ 17 Conclusion: The Solution of Planar QCD
Our second main outcome is dynamical: we demonstrated that the internal Majorana fermions (the “elves”) on the worldsheet provide the precise algebraic mechanism required for planar factorization at intersections. Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions. The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/96-1/96 and $47/96$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
One may view this construction as a geometric realization of the Master Field [54]—the existence of a purely classical description for the large-$N_{c}$ limit of QCD. While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, it finds an exact geometric representation in loop space. The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface. Consequently, the phase-space path integral over quark loops, once transformed into two-dimensional boundary twistor variables, exhibits exact geometric localization governed by Catastrophe Theory. The true Master Field of Planar QCD is thus a critical classical trajectory in twistor space—a geometric structure conceptually parallel to the spectral curves of Seiberg-Witten theory.
@@
Ultimately, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry. The extraction of the exact mass spectrum reduces from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem. By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition.
For the fundamental one-pole trajectory, this yields the unified parametric Regge trajectory (1.1) with trigonometric functions of the twistor branch point degree $a$, and $q=0$ for the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 for the shifted natural parity ($\rho$) trajectory. Although these functions are mathematically non-linear, they are numerically remarkably close to linear Regge trajectories. This analytic solution passes through the experimental data on Figure 15. The only outlier ($3\sigma$) is $K_{3}(2320)$, which is marked as “unconfirmed” in the PDG tables. The exact mass spectrum in our solution of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
@@ Acknowledgements
No data was created in this paper. The analytic computations of the Magnus expansion of the Momentum loop equations were performed in the Mathematica ® notebook published on the Wolfram Cloud[34]. The computations of the Regge trajectories and fitting the slopes to the meson mass data are done in Mathematica ® notebook published on the Wolfram Cloud[35].
@@ D.2 The Euclidean Metric and Parity Projection
To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$. The flat line element transforms as $|dz|^{2}=|z|^{2}(d\sigma_{E}^{2}+d\tau_{E}^{2})=e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$. Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
@@
$ds^{2}_{E}=\exp\left(2\rho\right)e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$
$=R^{2}\left(1+e^{2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
To evaluate the physical area, we analytically continue this metric into Minkowski spacetime via a double Wick rotation. The angular parameter $\tau_{E}$ serves as the target space time. We perform a Wick rotation on the spatial worldsheet coordinate into the physical light-cone regime:
@@
Under this transformation, the exponential strictly rotates into a complex phase: $e^{2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})+\mathrm{i}\sin(2\sigma_{M})$.
Because the macroscopic QCD string is strictly parity-conserving, the physical area is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations. The conjugate right-chiral surface identically provides the inverse complex phase $e^{-2\mathrm{i}\sigma_{M}}$. Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the area:
$\text{Re}\left[\exp\left(2\rho\right)e^{2\mathrm{i}\sigma_{M}}\right]$ $=R^{2}\big(1+\cos(2\sigma_{M})\big)$
$=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
@@ Appendix E The Pinch Singularity and the Topological Barrier
To mathematically illustrate the appearance of the conjugate pole at $1/\bar{z}_{0}$ and the resulting impenetrable topological barrier at the unit circle, consider the evaluation of the Liouville kinetic anomaly term in the effective action. In the case of a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
$I(z_{0})=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
In order to continue this integral as an analytic function of $z_{0}$ we rewrite the integrand as a total derivative with respect to $\bar{z}$:
$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}<1}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $z\bar{z}=1$:
$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
@@
$I(z_{0})=\frac{-\mathrm{i}}{2}\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
@@
At $z_{0}=0$ this integral reduces to
$I(0)=\frac{-\mathrm{i}}{2}\int_{0}^{2\pi}\mathrm{i}\,d\theta\left(-\mathrm{i}\theta\right)=\frac{-\mathrm{i}}{2}\frac{(2\pi)^{2}}{2}=-\mathrm{i}\pi^{2}$ (E.5)
which matches with the computation of the Minkowski-continued area in the previous section.
@@ Appendix F Algebraic Derivation of Spin Holonomy and Trace T
$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{2\pi i}&0\\ 0&e^{-2\pi i}\end{pmatrix}$ (F.4)
@@
The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$, which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\pi n$ depending on the parity of $L$.
Appendix G Geometric Potentials and Universal Trajectories
In this appendix, we decouple the unknown dimensionful string tension $\sigma$ from the geometric quantization conditions to derive the parametric functions $h(a)$ and $g(a)$, their singularities, and their roots.
We formulate the system directly in terms of the fractional winding parameter $a$. As derived in Section 15.5, the macroscopic spatial rotation angle of the string is exactly $\phi=4\pi a$. We isolate the string tension from the boundary potential $F(a)$ by defining the dimensionless geometric function:
$\tilde{F}(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (G.1)
The boundary potential is $F(a)=\frac{1}{\pi\sigma}\tilde{F}(a)$. Defining the physical mass-squared as $m^{2}=E^{2}=\sigma h(a)$, the unknown tension $\sigma$ factors to the numerator of the mass equation and cancels out of the angular momentum equation entirely.
The total geometric phase for the $q$-shifted trace sector is:
$\Phi_{q}(a)=-E^{2}F(a)+\frac{\pi a^{2}}{12}+4\pi aJ+2\pi qa$ (G.2)
where $q=0$ for the unshifted $\pi/K$ integral, and q=1q=-1 for the shifted $\rho$ and Scalar integrals.
We impose the saddle-point condition $\Phi_{q}^{\prime}(a)=0$ to isolate the angular momentum terms:
$4\pi aJ+2\pi qa=-\frac{\pi a^{2}}{6}+aE^{2}F^{\prime}(a)$ (G.3)
Substituting this into the Bohr-Sommerfeld quantization condition $\Phi_{q}(a)=2\pi n_{w}$ (where $n_{w}=0$ for the principal string sector), we obtain the algebraic relation:
$-\frac{\pi a^{2}}{12}-E^{2}F(a)+aE^{2}F^{\prime}(a)=0$ (G.4)
Using the differential identity $aF^{\prime}(a)-F(a)=\frac{1}{\pi\sigma}\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)$, we isolate $m^{2}=\sigma h(a)$ to obtain the mass function:
$h(a)=\frac{a^{2}\pi^{2}}{12\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)}$ (G.5)
Simplifying that expression, we get an explicit trigonometric answer:
$h(a)=-\frac{\pi^{2}(a-1)^{2}a^{2}\csc^{4}(\pi a)}{24(-2a+4\pi(a-1)a\cot(\pi a)+1)}$ (G.6)
To obtain the angular momentum function, we substitute $E^{2}=\sigma h(a)$ back into the saddle-point equation. Defining the target angular momentum as $u=J+q/2$, we obtain:
$u=\frac{a\left(a\tilde{F}^{\prime}(a)-2\tilde{F}(a)\right)}{48\left(\tilde{F}(a)-a\tilde{F}^{\prime}(a)\right)}$ (G.7)
This expression can be dramatically simplified. Dividing the numerator and denominator by $\tilde{F}(a)$ to isolate the logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)$, we obtain a partial fraction decomposition:
$\frac{\tilde{F}^{\prime}(a)}{\tilde{F}(a)}=\frac{2a+48u}{a(a+48u)}=\frac{1}{a}+\frac{1}{a+48u}$ (G.8)
Substituting the explicit logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)=4\pi\cot(\pi a)+\frac{1}{1-a}$ and combining the kinematic poles yields the exact trigonometric spectral equation:
$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (G.9)
The Regge trajectory can also be simplified to the following
$J+\frac{q}{2}=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{a}{48}$ (G.10)
G.1 Singularities and the Ground States
The geometry of the functions is dominated by a kinematic pole at the speed-of-light boundary, $a\rightarrow 1/2$. At this limit, the left-hand side of (G.9) strictly vanishes ($\cot(\pi/2)=0$). For the equation to hold, the right-hand side must also vanish, which dynamically forces the angular momentum to diverge ($u\rightarrow\infty$) and forms the linear macroscopic Regge asymptote. Evaluating the derivative $\tilde{F}^{\prime}(1/2)=8$ in the mass formula (G.5) restores the macroscopic Regge slope $J\approx 2m^{2}/(\pi^{2}\sigma)$. The asymptotic expansion at larger $s=m^{2}/\sigma$ corresponds to an inverted expansion around $a\to 1/2-\epsilon$. The expansion of trajectory goes as follows:
$J(s)$ $=\frac{2s}{\pi^{2}}+B_{0}+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+O(s^{-5});$ (G.11)
$B_{0}$ $=-\left(\frac{q}{2}+\frac{1}{96}\right);$ (G.12)
$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}$ (G.13)
$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}$ (G.14)
The coefficients $B_{1},B_{3}$ are numerically so small, that the trajectory is indistinguishable from linear for $J>1$.
When we compute the roots for the ground states, the explicit spectral equation (G.9) yields the mass eigenvalues:
  • 1.
    The Light Pseudoscalar ($q=0,u=0$): Solving (G.9) yields $a_{\pi}\approx 0.4268$. Substituting into $h(a)$ gives the mass ratio $m_{\pi}^{2}/\sigma\approx 0.0375$. The geometry produces a light ground state without symmetry breaking parameters.
  • 2.
    The Vector Ground State (q=1,u=1/2q=-1,u=1/2): The topological shift forces the string to spin faster. Solving (G.9) occurs at $a_{s}\approx 0.4987$, yielding $m_{s}^{2}/\sigma\approx 1.9782$. The geometry predicts the vector state to be massive.
The continuous generation of the parametric trajectories $u=J+q/2$, the numerical roots, and the data fits to $\sigma$ are performed using the companion Mathematica notebook [35].
G.2 The Endpoint Singularity and the Fold Catastrophe
While the large-$s$ asymptotic expansion (G.11) governs the strictly linear macroscopic Regge regime, the exact parametric geometry of the twistor string dictates a hard geometric boundary at small $s$. This boundary manifests not as a smooth continuation toward $s\to$-$\infty$, but as a fold catastrophe in the mapping between the spectral parameter $a$ and the physical observables.
To expose this geometric termination, we rewrite the exact parametric equations for the mass ratio $s(a)=h(a)=m^{2}/\sigma$ and the target angular momentum $u(a)=J+q/2$. From (G.6) and (G.8), we define the auxiliary functions:
$\mathcal{P}(a)$ $=\frac{\pi^{2}a(1-a)}{24\sin^{4}(\pi a)},$ (G.15)
$\mathcal{R}(a)$ $=4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}.$ (G.16)
The spectral equations then take the remarkably compact algebraic form:
$s(a)$ $=\mathcal{P}(a)\mathcal{R}(a)^{-1},$ (G.17)
$48u(a)$ $=\mathcal{R}(a)^{-1}-a.$ (G.18)
A fundamental property of the Rigid Twistor String is the existence of a simultaneous mathematical endpoint $a_{c}$ where the first derivatives of both the mass-squared and the angular momentum vanish exactly. We can rigorously prove this simultaneous singularity by differentiating the target angular momentum:
$48u^{\prime}(a)$ $=-\mathcal{R}^{\prime}(a)\mathcal{R}(a)^{-2}-1=0$
$\implies-\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}.$ (G.19)
Next, we differentiate the mass function and multiply by $\mathcal{R}(a)^{2}$:
$s^{\prime}(a)\mathcal{R}(a)^{2}=\mathcal{R}(a)\left[\mathcal{P}^{\prime}(a)+\mathcal{P}(a)\mathcal{R}(a)\right].$ (G.20)
By explicitly taking the logarithmic derivative of $\mathcal{P}(a)$, we uncover an exact geometric identity:
$\frac{\mathcal{P}^{\prime}(a)}{\mathcal{P}(a)}=\frac{1-2a}{a(1-a)}-4\pi\cot(\pi a)\equiv-\mathcal{R}(a).$ (G.21)
Consequently, the term in the brackets vanishes identically. Thus, we arrive at the exact algebraic statement: $s^{\prime}(a_{c})=0$ precisely whenever $u^{\prime}(a_{c})=0$.
In the physical $(s,J)$ plane, this simultaneous extremum signifies a fold catastrophe. The parametric Regge trajectory reaches a hard geometric endpoint where the physical slope $dJ/ds=u^{\prime}(a)/s^{\prime}(a)$ resolves (via L’Hôpital’s rule) to a finite limit $S_{0}$, while the trajectory curvature diverges to infinity, $J^{\prime\prime}(s)\to\infty$.
The critical parameter $a_{c}$ is determined by the exact transcendental equation $\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}$, which evaluates to:
$4\pi^{2}\csc^{2}(\pi a)-\frac{1-2a+2a^{2}}{a^{2}(1-a)^{2}}$
$=\left(4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}\right)^{2}.$ (G.22)
Solving this within the fundamental domain $a\in(0,1/2)$ yields the universal critical parameters:
Crit. Param. ($a_{c}$): $\approx 0.311277$
Thres. Mass Rat. ($s_{c}=m_{c}^{2}/\sigma$): $\approx 0.027791$
Crit. $\pi$ Spin ($q=0$): $J_{c}\approx-0.003378$
Crit. $\rho$ Spin (q=1q=-1): $J_{c}\approx 0.496622$ (G.23)
Figure 16: The exact geometric Regge trajectories plotted down to the mathematical endpoint $(s_{c},J_{c})$. The trajectory departs from this boundary with a finite slope ($S_{0}\approx 0.1118$) that is smaller than the asymptotic slope at large $s$ ($\approx 2/\pi^{2}\approx 0.2026$). The infinite curvature $J^{\prime\prime}(s)\to\infty$ at the exact endpoint is visually imperceptible. To the eye, this looks like a perfectly linear trajectory ending abruptly at this point.
At this mathematical endpoint, the Regge trajectory departs with a finite slope $S_{0}\approx 0.1118$ and infinite curvature (see Fig. 16). While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15). The physical branch of the spectrum strictly corresponds to $u\geq 0$. As established previously, the physical ground state terminates at $a_{min}$ where $u=0$. For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
Because the physical root $a_{min}$ sits comfortably above the critical parameter $a_{c}$ ($0.4268>0.3113$), the trajectory mathematically terminates well before the string can experience the fold catastrophe. This mass gap ($s(a_{min})>s_{c}>0$) acts as an impenetrable geometric shield. It guarantees that the Regge trajectory never crosses into the $s<0$ regime, rigorously protecting the Planar QCD spectrum from tachyonic instabilities while enforcing the exact reality of all physical eigenvalues. The phenomenological Regge “intercepts” $\alpha_{\pi,\rho}(0)$ are thus not true physical states, but simply linear extrapolations of the large-$s$ asymptotic expansion bridging over this hidden, non-linear geometric boundary.
While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15). The physical branch of the spectrum strictly corresponds to $u\geq 0$. As established previously, the physical ground state terminates at $a_{min}$ where $u=0$. For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
Appendix H Irrationality of the Physical Branch Points
As discussed in Section 15.8, rational values of the branch point degree $a=m/n$ would cause the twistor trajectory to close after finitely many windings. In that case, all distinct states would be counted with just $2n$ windings, and our mechanism for the Exact WKB spectrum would disappear. In this appendix, we rigorously prove that the physical spectrum is mathematically protected from these rational degeneracies.
Theorem. For any rational target angular momentum $u=J+q/2\geq 0$, the exact spectral equation (G.9) possesses no rational solutions for $a\in(0,1/2)$.
Proof. The exact geometric quantization is governed by the spectral equation derived in Appendix G:
$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+48u}$ (H.1)
Assume, for the sake of contradiction, that the physical branch point $a$ is a rational number $m/n\in(0,1/2)$. Since the target momentum $u$ is also rational by definition, the algebraic expression on the right-hand side evaluates to a purely rational number.
On the left-hand side, it is a foundational property of trigonometric functions that for any rational argument $a$, the value of $\cot(\pi a)$ is an algebraic number. Furthermore, for $a\in(0,1/2)$, the value $\cot(\pi a)$ is strictly positive and therefore non-zero. By the Lindemann–Weierstrass theorem, $\pi$ is a transcendental number. The product of a transcendental number and a non-zero algebraic number is strictly transcendental.
Consequently, the left-hand side is a transcendental number, while the right-hand side is a rational number. This is a strict mathematical contradiction. Therefore, the branch point parameter $a$ cannot be a rational number. $\blacksquare$